OpenStudy (anonymous):

Simplify the following expression. 2x − 6y + 3x2 + 7y − 14x

OpenStudy (anonymous):

2x-6y+3x^2+7y-14x = 3x^2+(2-14)x+(7-6)y= 3x^2-12x=y